# FastSet: Parallel Claim Settlement

### v1

@@ 4.14 Time Stamping
An alternative is to have time service providers on the FastSet network, whose job is to verify time claims made by the clients. Specifically, clients can send messages along with the timestamp they claim for the message, and have time verifiers sign the message and verify the claimed timestamp, in Step 2 of FastSet (see Figure 1). Of course, the verifiers would need to allow an error margin to compensate for network delays. [^1] A national network (WAN) could see delays between 30 and 50 ms, and international connections (WWAN) might have delays between 100 and 300 ms. This can be problematic if other verifiers are also needed in the same message. Variations of FastSet may choose to allow different sets of verifiers for different tasks, each with its own quorum.
Considering that perfect timestamps are off the table and that validators will likely be reasonably good timekeepers anyway, implementations of FastSet may opt for having their validators also validate timestamps. For all claims, not only those including explicit timestamps. Specifically, suppose that the messages sent by clients in Step 1 of FastSet have the form $\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}^{\tau}$, that is, the timestamp $\tau$ is part of the message, same like the nonce $n$, and not a special claim. Then each validator checks $\tau$ as well in Step 3, within an adequate error bound to allow for network delays. Allowing clients to timestamp their messages has the additional benefit that messages that are late can be discarded at any moment by proxies and validators which may gossip, and, with some care, validators may also clear their $v(a).pending$ when it contains long expired messages. Finally, this also allows the contract language to have a time construct, which is interpreted by the validators as the time given by $\tau$.
@@
Here we assumed that claim blocks are executed atomically, that is, different block claims accessing the same shared variable will not interleave their claims. This is important, for example, in the constructor above where the shared variables time can be read and written by various instances at the same time. The overall effect on the shared variables will be the same once all instances are settled regardless of the order in which the instances are processed, but it is important that only one instance at a time accesses the shared variables. We are aware that the atomicity of claim blocks is a very strong assumptions and plan as future work to weaken it by adding synchronization objects (mutexes).
@@ 4.15 Auctions
The auctioneer creates the contract above, at the same time sending it the ITEM for bidding as well as a BIDDING_TIME for how long bidding is allowed. For example, $AliceTicket|=\langle\verb|contract(AUCTION[ticket_item,1000])|\rangle_{\texttt{alice}}$. For simplicity we assume only one item, which is a token as in Section 4.12. This item is locked in the contract until the auction ends. The contract initializes the highestBidder as the contract owner, alice in our case, so the contract owner can redeem the item in case nobody bids on it during the specified period.
@@
At any moment during or after the auction ends, the bidders can withdraw any available funds from their instances, e.g.:
@@
Suppose now that half of the validators receive/process the two bids in one sequence in their Step 7, say bob first and charlie second, while the other half in the other sequence, charlie first and bob second. In both cases the contract will hold 30 set, charlie will be the highest bidder, bob’s instance balance of 100 set is unaffected, and charlie’s instance balance is 20 set. However, in the second case, bob’s bid never took place, in the sense that the body of the conditional statement was not executed. But, importantly, bob’s bid block was still valid, so in the end all validators are in the same state and with the same claims settled. Since bob’s instance balance was not affected on any of the validators, his next auction3.bid(40)bobsubscriptdelimited-⟨⟩monospace-$auction3.bid(40)|\rangle_{\texttt{bob}}$ correctly outbids charile on all validators. We leave it as an exercise to the curious reader to notice that weak independence would be violated if we replaced the conditional statement with a guard that invalidated the block when the amount was not higher than the highest bid. In Section 4.17 we discuss this case in more depth and propose an extension of FastSet that would work with such examples where weak independence is allowed to be temporarily broken. We also leave it as an exercise to the reader to notice that the weak independence assumption would also be violated if the previously highest bid would be returned directly to the owner of the instance, instead of to the instance itself. We discuss this case in more depth in Section 4.17 as well.
@@ 4.16 AppChains and Blockchains
For concreteness, let’s assume the verifiable compute context discussed in Section 4.8, where EVM is settled as a language claim and VERIFIED_COMPUTE is a contract that allows its instances to make use of a set of specialized verifiers to verify compute claims using EVM and then settle them as claims of the form:
@@
Above, alice submitted a transaction on the chain, alice_trans, and 1 set to the contract. This could be, for example, a bridge transfer, which mints 1 set token in alice’s account on evmchain. We are not concerned with the integrity or semantics of the specific chain transactions here. These can be verified by the chain contract itself, or by its verifiers, or can be generated to be semantically-correct-by-construction using variants of submit blocks, e.g., evmchain_alice.bridge(1)alicesubscriptdelimited-⟨⟩monospace-$evmchain_alice.bridge(1)|\rangle_{\texttt{alice}}$ (for brevity, we do not define bridge here).
@@ 4.17 Beyond Weak Independence
First, consider the apparently innocent change in the AUCTION contract in Section 4.15, where the new highest bidder pays back the previously highest bidder owner, instead of their instance. That is, replace instance with instance.owner in the bid block (and remove .owner in the end block). Consider the same scenario discussed in Section 4.15, where bob and charlie submit their bids concurrently, and half the validators process bob first and then charlie, while the other half charlie first and bob second. In both cases the contract will hold 30 set, charlie will be the highest bidder, and charlie’s instance balance is 20 set. However, in the first case bob’s balance is 25 set more while bob’s instance balance is 75 set, while in the second case bob’s balance is unaffected while bob’s instance balance is 100 set. It may look like there is no problem, because bob’s total balance in both its account and in its instance stays the same, and he always has the option to withdraw from the instance. However, bob effectively lost 25 set, because he will only be able to get validator quorum on withdrawals of up to 75 set. Additionally, from here on, bob’s account will show different balances on different validators, without any chance for the validators to ever converge on bob’s balance.
Let us next consider another change to the AUCTION contract in Section 4.15, where the conditional statement in bid is replaced with its condition as a guard followed by the same commands. That is, bid is invalid now when the amount is insufficient, instead of silently valid. Now the case when bob’s bid is processed after charlie’s is invalid, thus breaking the weak independence assumption. As a consequence, bob’s bid cannot be settled. In fact, it has no chance to ever be settled for that half of the validators, regardless of how the auction proceeds, not even after the auction end. So bob’s bid will stay forever in $v(\verb|auction3|).\textit{pending}$ for all those validators. Not only this is garbage both computationally and space-wise for half the validators, but also it locks the auction3 account at the current nonce on those validators, which means that auction3 is effectively forever locked together with bob’s 100 set, because it will never be able to achieve quorum on any message anymore.
@@
setl’s monotonicity enforcement resolution makes the modified auction contract work as intended: fast, correct, minimal on resources. Indeed, each bidder submits their bids unrestricted by anything but the deadline, as it should be, without any sequencing or waiting required, as it should be, and only the contract needs to store data, and only two integers: the highest bidder and its amount. Correctness means that from all validators’ perspectives, the highest bidder gets the item from the auctioneer and all other bidders get their money back. It follows easily by induction. Suppose that the property holds for $n$ bidders, and consider one more bid. Let $v$ be any validator. If the new bid is higher than the previously highest bid for $v$, then the bid claim block will set the new bid as the highest and the previously highest bidder will be transferred its money back by the new bidder (which can also be the same bidder). So the inductive property holds in this case. If the new bid happens to be lower than the previously highest bid for $v$, which can happen because of the receive and settlement order in Step 7 of the FastSet protocol on $v$, then the bid claim block will be invalid and therefore, by the monotonicity enforcement resolution above, will not change the state of $v$ (but it will increase the nonce of the bidder, so it can continue to participate in bidding). So the property holds in this case, too.
@@ 5 Conclusion
This paper introduced FastSet, a replica-based distributed protocol for embarrassingly parallel settlement of claims. A claim is any statement that comes with a proof, such as a payment or more generally a digital asset transfer, a realized blockchain transaction or more transactions that form a block, an execution of a program in a programming language or virtual machine, an AI model inference or fine tuning, a TEE execution, a vote, an auction bid, among many others. A proof is anything that is verifiable and is accepted by the application or user/account that uses the claim: a signature (simple, aggregated, TEE, etc), a cryptographic/ZK proof, a mathematical proof, a formal semantics derivation, a program (re)execution, and so on. The verification of all proofs which are not signatures is deferred to special service providers, called verifiers which are account holders like any other applications/users. The replicas, called validators, only validate signatures, regarded as the simplest and fastest to check proofs, and settle the claims. Each claim can be verified, validated, and settled optimally: independently and in parallel with any other claim. The validators need not communicate with each other, so there is no consensus in the strict sense of the word as used in blockchains. Specifically, FastSet is not strongly consistent [12], but it is strongly eventually consistent [21].
FastSet drew inspiration from three relatively recent works, published after 2019, which in our view have brought irrefutable evidence that strong consistency in general, and blockchains in particular, are not necessary in order to prevent the double spending attack problem: the consensus number of a cryptocurrency [9], FastPay [3], and POD [2]. These works demonstrated that the peer-to-peer decentralized payment system challenge that was notoriously solved by Bitcoin [16] and has stimulated more than a decade of sustained research, engineering and investments advancing the blockchain technology to what it is today, can also be solved differently. Quite completely differently, in fact, with no chains (or total order) and no blocks, and without any theoretical limitations on the number of transactions per second. FastSet brings a general mechanism to settle any claims that are verifiably true, including universal programmability: any computations done using any programming models, languages or VMs. At a minimum, FastSet is to FastPay what Ethereum is to Bitcoin.
@@
FastSet can be extended or optimized in several directions. In the current design, FastSet validators use and maintain their own local state and need not communicate with each other. Consistency is maintained thanks to weak independence of claims and commutativity of claim settlement. However, broadcasting to a large number of validators may not be practical in some situations, so gossip communication among the validators can be allowed to more efficiently broadcast new claims across the network, using related technologies in coding theory such as Optimum [5] (https://getoptimum.xyz) and set reconciliation [23]. We currently assume that all validators are involved in all claim validations, which means that for any claim settlement, a quorum of the entire base of validators must be reached. Adding more validators to the network, in theory, will not increasing the latency of FastSet because validators work independently from each other. However, it will increase the network traffic. A possible solution is to adopt a committee-based approach like in Algorand [7] and others to enable even more parallel settlement and thus further improve performance. In the current design, a certified but (temporarily) invalid claim is marked as pre-settled and then queued in Step 6. In a worst and most unfortunately scenario, the queue can grow indefinitely. For example, a possibly malicious application may DOS attack some validators by delaying them from seeing messages which would turn the invalid certified claim valid. There exist many possible solutions. A time-based solution, based on the timestamping feature introduced in Section 4.14, may allow validators to discard invalid certified claims that have become outdated and thus allow FastSet to offer a perfect-past consensus result like in POD [2]. Another possible solution is to enable the processing and settlement of invalid claims, generalizing FastPay [3], where invalid payments can still be processed and settled, resulting in (temporary) negative account balances. We leave these extensions and optimizations as future work.
@@ A.1 Two Types of Accounts
setl’s user-driven accounts and contract-based accounts are similar to Ethereum’s externally-owned and, resp., contract accounts. Both user-driven and contract-based accounts can receive, hold, and send native tokens and interact with deployed contracts. Their key differences are also similar to those between Ethereum’s externally-owned accounts and contract accounts, listed as follows:
@@
    • Creating an account costs nothing;
@@
    • Controlled by a pair of public and private keys.
@@
    • Creating an account has a cost because it is using network storage;
@@ A.2 Contracts and Instances
Both contracts and instances are setl accounts, and they both are contract-based accounts, meaning that their behaviors are predetermined by setl scripts. Their account addresses are their unique IDs, which are publicly known and thus can be used as references to them. Contracts and instances both have fields and blocks. Fields can hold typed values and they constitute the account states, while blocks are prescribed scripts associated to the contract/instance accounts. This way, setl is similar to an object-orient (OO) programming language, where contracts and instances are similar to classes and objects, while fields and blocks are similar to class members (fields and methods). However, as we will shortly see, setl’s contracts, instances, fields, and blocks differ from their counterparts in OO languages in several important and fundamental ways.
@@
Via contracts and instances, setl imposes a principled scheme for various parties and stakeholders to concurrently cooperate and coordinate in a shared time-space. In this scheme, a contract creator sets the stage, by specifying the fields and blocks of the contract, as well as the fields and blocks of its instances. From then on, the stage has been set. The contract account can only behave according to the prescribed contract blocks at the request of the contract creator, while every instance account can only behave according to the prescribed instance block at the request of the instance creator.

### v3

@@ 4.14 Time Stamping
An alternative is to have time service providers on the FastSet network, whose job is to verify time claims made by the clients. Specifically, clients can send messages along with the timestamp they claim for the message and have time verifiers sign the message and verify the claimed timestamp (Step 2 of FastSet in Figure 1). Of course, the verifiers would need to allow an error margin to compensate for network delays. [^2] A national network (WAN) could see delays between 30 and 50 ms, and international connections (WWAN) might have delays between 100 and 300 ms. This can be problematic if other verifiers are also needed in the same message. Variations of FastSet may choose to allow different sets of verifiers for different tasks, each with its own quorum.
Considering that perfect timestamps are off the table and that validators will likely be reasonably good timekeepers anyway, implementations of FastSet may opt for having their validators also validate timestamps—for all claims, not only those including explicit timestamps. Specifically, suppose that the messages sent by clients in Step 1 of FastSet have the form $\langle c_{1}\,c_{2}\,...\,c_{k},n\rangle_{a}^{\tau}$, that is, the timestamp $\tau$ is part of the message, just as the nonce $n$, and not a special claim. Then each validator also checks $\tau$ in Step 3, within an adequate error bound to allow for network delays. Allowing clients to timestamp their messages has the additional benefit that messages that are late can be discarded at any moment by proxies and validators which may gossip, and, with some care, validators may also clear their $v(a).pending$ when it contains long expired messages. Finally, this also allows the contract language to have a time construct, which is interpreted by the validators as the time given by $\tau$.
@@
Here we assumed that claim blocks are executed atomically, that is, different block claims accessing the same shared variable will not interleave their claims. This is important, for example, in the constructor above where the shared variables time can be read and written by various instances at the same time. The overall effect on the shared variables will be the same once all instances are settled regardless of the order in which the instances are processed, but it is important that only one instance at a time accesses the shared variables. We are aware that the atomicity of claim blocks is a very strong assumption and plan as future work to weaken it by adding synchronization objects (mutexes).
@@ 4.15 Auctions
The auctioneer creates the contract above, at the same time sending it the ITEM for bidding, as well as a BIDDING_TIME for how long bidding is allowed, e.g., $AliceTicket|=\langle\verb|contract(AUCTION[ticket_item,1000])|\rangle_{\texttt{alice}}$. For simplicity we assume only one item, which is a token as in Section 4.12. This item is locked in the contract until the auction ends. The contract initializes the highestBidder as the contract owner, alice in our case, so the contract owner can redeem the item in case nobody bids on it during the specified period.
@@
At any moment during or after the auction ends, the bidders can withdraw any available funds from their instances, e.g.,
@@
Suppose now that half of the validators receive/process the two bids in one sequence in their Step 7, say bob first and charlie second, while the other half in the other sequence, charlie first and bob second. In both cases the contract will hold 30 tokens, charlie will be the highest bidder, bob’s instance balance of 100 tokens is unaffected, and charlie’s instance balance is 20 tokens. However, in the second case, bob’s bid never took place, in the sense that the body of the conditional statement was not executed. But, importantly, bob’s bid block was still valid, so in the end all validators are in the same state and with the same claims settled. Since bob’s instance balance was not affected on any of the validators, his next auction3.bid(40)bobsubscriptdelimited-⟨⟩monospace-$auction3.bid(40)|\rangle_{\texttt{bob}}$ correctly outbids charile on all validators. We leave it as an exercise to the curious reader to notice that weak independence would be violated if we replaced the conditional statement with a guard that invalidated the block when the amount was not higher than the highest bid. In Section 4.17 we discuss this case in more depth and propose an extension of FastSet that would work with such examples where weak independence is allowed to be temporarily broken. We also leave it as an exercise to the reader to notice that the weak independence assumption would also be violated if the previously highest bid would be returned directly to the owner of the instance, instead of to the instance itself. We discuss this case in more depth in Section 4.17 as well.
@@ 4.16 AppChains and Blockchains
For concreteness, let us assume the verifiable compute context discussed in Section 4.8, where EVM is settled as a language claim and VERIFIED_COMPUTE is a contract that allows its instances to make use of a set of specialized verifiers to verify compute claims using EVM and then settle them as claims of the form:
@@
Above, alice submitted a transaction on the chain, alice_trans, and 1 token to the contract. This could be, for example, a bridge transfer, which mints 1 token in alice’s account on evmchain. We are not concerned with the integrity or semantics of the specific chain transactions here. These can be verified by the chain contract itself, or by its verifiers, or can be generated to be semantically-correct-by-construction using variants of submit blocks, e.g., evmchain_alice.bridge(1)alicesubscriptdelimited-⟨⟩monospace-$evmchain_alice.bridge(1)|\rangle_{\texttt{alice}}$ (for brevity, we do not define bridge here).
@@ 4.17 Beyond Weak Independence
First, consider the apparently innocent change in the AUCTION contract in Section 4.15, where the new highest bidder pays back the previously highest bidder owner, instead of their instance. That is, replace instance with instance.owner in the bid block (and remove .owner in the end block). Consider the same scenario discussed in Section 4.15, where bob and charlie submit their bids concurrently, and half the validators process bob first and then charlie, while the other half charlie first and bob second. In both cases the contract will hold 30 tokens, charlie will be the highest bidder, and charlie’s instance balance is 20 tokens. However, in the first case bob’s balance is 25 tokens more while bob’s instance balance is 75 tokens, while in the second case bob’s balance is unaffected while bob’s instance balance is 100 tokens. It may look like there is no problem, because bob’s total balance in both its account and in its instance stays the same, and he always has the option to withdraw from the instance. However, bob effectively lost 25 tokens, because he will only be able to get validator quorum on withdrawals of up to 75 tokens. Additionally, from here on, bob’s account will show different balances on different validators, without any chance for the validators to ever converge on bob’s balance.
Let us next consider another change to the AUCTION contract in Section 4.15, where the conditional statement in bid is replaced with its condition as a guard followed by the same commands. That is, bid is invalid now when the amount is insufficient, instead of silently valid. Now the case when bob’s bid is processed after charlie’s is invalid, thus breaking the weak independence assumption. As a consequence, bob’s bid cannot be settled. In fact, it has no chance to ever be settled for that half of the validators, regardless of how the auction proceeds, not even after the auction ends. So bob’s bid will stay forever in $v(\verb|auction3|).\textit{pending}$ for all those validators. Not only is this garbage both computationally and space-wise for half the validators, but also it locks the auction3 account at the current nonce on those validators, which means that auction3 is effectively forever locked together with bob’s 100 tokens, because it will never be able to achieve quorum on any message anymore.
@@
setl’s monotonicity enforcement resolution makes the modified auction contract work as intended: fast, correct, and minimal on resources. Indeed, each bidder submits their bids unrestricted by anything but the deadline, as it should be, without any sequencing or waiting required, as it should be, and only the contract needs to store data, and only two integers: the highest bidder and its amount. Correctness means that from all validators’ perspectives, the highest bidder gets the item from the auctioneer and all other bidders get their money back. It follows easily by induction. Suppose that the property holds for $n$ bidders, and consider one more bid. Let $v$ be any validator. If the new bid is higher than the previously highest bid for $v$, then the bid claim block will set the new bid as the highest and the previously highest bidder will be transferred its money back by the new bidder (which can also be the same bidder). So the inductive property holds in this case. If the new bid happens to be lower than the previously highest bid for $v$, which can happen because of the receive and settlement order in Step 7 of the FastSet protocol on $v$, then the bid claim block will be invalid and, therefore, by the monotonicity enforcement resolution above, will not change the state of $v$ (but it will increase the nonce of the bidder, so it can continue to participate in bidding). So the property holds in this case, too.
@@ 5 Conclusion
This paper introduced FastSet, a replica-based distributed protocol for embarrassingly parallel settlement of claims. A claim is any statement that comes with a proof, such as a payment or more generally a digital asset transfer, a realized blockchain transaction or more transactions that form a block, an execution of a program in a programming language or virtual machine, an AI model inference or fine tuning, a TEE execution, a vote, an auction bid, among many others. A proof is anything that is verifiable and is accepted by the application or user/account that uses the claim: a signature (simple, aggregated, TEE, etc), a cryptographic/ZK proof, a mathematical proof, a formal semantics derivation, a program (re)execution, and so on. The verification of all proofs which are not signatures is deferred to special service providers, called verifiers which are account holders like any other applications/users. The replicas, called validators, only validate signatures, regarded as the simplest and fastest way to check proofs, and settle the claims. Each claim can be verified, validated, and settled optimally: independently and in parallel with any other claim. The validators need not communicate with each other, so there is no consensus in the strict sense of the word as used in blockchains. Specifically, FastSet is not strongly consistent [12], but it is strongly eventually consistent [21].
FastSet drew inspiration from three relatively recent works, published after 2019, which in our view have brought irrefutable evidence that strong consistency in general, and blockchains in particular, are not necessary in order to prevent the double spending attack problem: the consensus number of a cryptocurrency [9], FastPay [3], and POD [2]. These works demonstrated that the peer-to-peer decentralized payment system challenge that was notoriously solved by Bitcoin [16] and has stimulated more than a decade of sustained research, engineering and investments advancing the blockchain technology to what it is today, can also be solved differently. Quite completely differently, in fact, with no chains (or total order) and no blocks, and without any theoretical limitations on the number of transactions per second. FastSet brings a general mechanism to settle any claims that are verifiably true, including universal programmability: any computations done using any programming models, languages, or VMs. At a minimum, FastSet is to FastPay what Ethereum is to Bitcoin.
@@
FastSet can be extended or optimized in several directions. In the current design, FastSet validators use and maintain their own local state and need not communicate with each other. Consistency is maintained thanks to weak independence of claims and commutativity of claim settlement. However, broadcasting to a large number of validators may not be practical in some situations, so gossip communication among the validators can be allowed to more efficiently broadcast new claims across the network, using related technologies in coding theory such as Optimum [5] (https://getoptimum.xyz) and set reconciliation [23]. We currently assume that all validators are involved in all claim validations, which means that for any claim settlement, a quorum of the entire base of validators must be reached. Adding more validators to the network, in theory, will not increase the latency of FastSet because validators work independently from each other. However, it will increase the network traffic. A possible solution is to adopt a committee-based approach like the one in Algorand [7] to enable even more parallel settlement and thus further improve performance. In the current design, a certified but (temporarily) invalid claim is marked as pre-settled and then queued in Step 6. In the worst and most unfortunate scenario, the queue can grow indefinitely. For example, a possibly malicious application may DOS attack some validators by delaying them from seeing messages which would turn the currently-invalid certified claim valid. There exist many possible solutions. A time-based solution relying on the timestamping feature introduced in Section 4.14 may allow validators to discard invalid certified claims that have become outdated and thus allow FastSet to offer a perfect-past consensus result like in POD [2]. Another possible solution is to enable the processing and settlement of invalid claims, generalizing FastPay [3], where invalid payments can still be processed and settled, resulting in (temporary) negative account balances. We leave these extensions and optimizations as future work.
@@ A.1 Two Types of Accounts
setl’s user-driven accounts and contract-based accounts are similar to Ethereum’s externally-owned and, respectively, contract accounts. Both user-driven and contract-based accounts can receive, hold, and send native tokens and interact with deployed contracts. Their key differences are also similar to those between Ethereum’s externally-owned accounts and contract accounts, listed as follows:
@@
    • Cost nothing to create;
@@
    • Are controlled by a pair of public and private keys.
@@
    • Incur a creation cost because they are using network storage;
@@ A.2 Contracts and Instances
Both contracts and instances are setl accounts and both are contract-based accounts, meaning that their behaviors are predetermined by setl scripts. Their account addresses are their unique IDs, which are publicly known and thus can be used as references to them. Contracts and instances both have fields and blocks. Fields can hold typed values and they constitute the account states, while blocks are prescribed scripts associated to the contract/instance accounts. This way, setl is similar to an object-oriented (OO) programming language, where contracts and instances are similar to classes and objects, while fields and blocks are similar to class members (fields and methods). However, as we will shortly see, setl’s contracts, instances, fields, and blocks differ from their counterparts in OO languages in several important and fundamental ways.
@@
Via contracts and instances, setl imposes a principled scheme for various parties and stakeholders to concurrently cooperate and coordinate in a shared time-space. In this scheme, a contract creator sets the stage by specifying the fields and blocks of the contract, as well as the fields and blocks of its instances. From then on, the stage has been set. The contract account can only behave according to the prescribed contract blocks at the request of the contract creator, while every instance account can only behave according to the prescribed instance block at the request of the instance creator.
A contract creation by a user-driven account effectively creates an actor. The semantics of setl only allows the contract owner to send messages to the contract actor. Indeed, in setl, instance fields form the state of the contract instance actors, and these instance actors can only initiate instance blocks according to the contract. The fields which do not have an instance modifier form the state of the contract actor, and the blocks which do not have an instance modifier can only be generated by the contract owner. In setl, anybody can create instances of the contract, and each instance is an actor controlled by the its owner user-driven account. This way, the semantics of setl is faithful to the actor model in that actors only communicate with actors they have knowledge about.
A clarification is needed with regards to the fact that owners of contracts and of instances, respectively, can sign on their behalf. In Section 2, FastSet stipulates that each account $a$ can only submit claim blocks which are signed by $a$. In setl, on the other hand, each contract creation and each instance of it becomes an actor, which has its own state and address. However, it would be rather inconvenient and likely insecure and unmanageable by users to allow each actor to own its own private key to sign all its claim blocks. Instead, setl allows the contract-based actors to only be controled by their owners, technically just other accounts that have a private key and thus can sign messages. That is, the unique owner of the created contract, or the unique owner of some instance of it, is the only entity that can sign blocks on behalf of the address corresponding to the associated contract or instance actor. This is not a modification of the FastSet protocol, but rather a particular implementation of it. Indeed, a single private key can be used to generate infinitely many public keys through various methods, like hierarchical deterministic (HD) wallets or different cryptographic algorithms. These different public keys are deterministically derived from the same private key and are used as contract-based actor addresses in setl, although there is only one initial pairing between a private key and its primary public key.